# A Geometric Substructure for Quantum Dynamics

### v4

@@ Abstract
The description of a closed quantum system is extended with the identification of an underlying substructure enabling an expanded formulation of dynamics in the Heisenberg picture. Between measurements a “state point” moves in an underlying multi-dimensional complex projective space with constant velocity determined by the quantum state vector. Following a measurement the point changes direction and moves with new constant velocity along one of several possible new orthogonal paths with probabilities determined by Born’s Rule. From this previously hidden substructure a new picture of quantum dynamics and quantum measurements emerges without violating existing no-go theorems regarding hidden variables. A natural generalisation to a Riemannian substructure is proposed, which suggests an interaction of quantum measurements with the background gravitational field.
email: ajb@maths.uq.edu.au Keywords: quantum dynamics; geometrical substructure; quantum measurements; quantum mechanics and gravitation.
ORCID: $0$-0001-5181-4483
@@ 1 Introduction
The description of the measurement process and the associated Born interpretation of the state vector have been the most contentious features of quantum mechanics since its inception. In the case of a conservative system the strangeness of the orthodox description is seen most clearly in the Heisenberg picture of quantum dynamics [8, 13, 14]. There the state vector $|\psi\rangle$ is a constant unit vector in a Hilbert space $\cal{H}$, possibly infinite-dimensional, between measurements at times $t_{0}$ and $t_{1}>t_{0}$, while self-adjoint operators ${\widehat{A}}(t)$, ${\widehat{B}}(t)$, $\dots$ representing observables evolve in time in accordance with Heisenberg’s equation of motion
@@
Here ${\widehat{H}}$ is the Hamiltonian operator. (For simplicity we consider only closed systems for which ${\widehat{H}}$ is time-independent.)
But in another and very different type of dynamical process that is assumed to reflect the interaction of the system with the measuring apparatus and the observer, a maximal set of commuting observables is chosen for each of the measurements and in each case the state vector is observed to move into a common eigenvector of that set in a non-deterministic way with a probability determined by Born’s Rule [8, 13, 14]. (Again for simplicity we ignore complications that arise if partially continuous spectra are involved.)
Suppose that at time $t_{0}$ a measurement is made of one such set of commuting observables and $|\psi\rangle$ is observed to move into one of their common eigenvectors. The state vector then remains constant for $t_{0}<t<t_{1}$. We may expand it during this time interval in terms of some chosen reference basis, a complete set of orthonormal vectors $|\varphi_{i}\rangle$, $i=1\,,2\,\dots\,$, possibly infinite in number, with constant complex expansion coefficients $\alpha^{i}$. Then
@@
Now consider a second, in general different maximal set of commuting observables ${\widehat{A}}(t)$, ${\widehat{B}}(t)$, $\dots$ evolving in time for $t>t_{0}$ in accordance with (1) and having a complete orthormal set of common eigenvectors $|\chi_{K}(t)\rangle$ for $K=1\,,2\,,\dots$ with components $\beta_{K}^{\,\,\,i}(t)$ in the same reference basis used for $|\psi\rangle$, so
@@
(In the special case when ${\widehat{H}}$ is one of the chosen set then all the observables in the set are constants of the motion and their common eigenvectors $|\chi_{K}\rangle$ and all their components $\beta_{K}^{\,\,\,i}$ are constant.)
@@
${\widehat{B}}(t_{1})\,|\psi\rangle={\widehat{B}}(t_{1})|\chi_{K^{\prime}}(t_{1})\rangle=b|\chi_{K^{\prime}}(t_{1})\rangle=b\,|\psi\rangle\,,\quad\dots$ (8)
@@ 2 Identifying a substructure
Our starting point is the observation that this dynamical process is strongly reminiscent of the behaviour of a free particle travelling with constant velocity between impulsive forces applied at $t_{0}$ and $t_{1}$. This suggests the association of the state vector $|\psi\rangle$ with the velocity vector of a point moving in a hitherto unidentified underlying space. Accordingly we identify $V\,\alpha^{i}$ with (the components of) the constant velocity vector $v^{i}$ of a “state point” moving in an underlying space $S$ say, with complex coordinates $z^{i}$ so that
$v^{i}=\frac{dz^{i}}{dt}\,,\qquad i=1\,,2\,,\,\dots$ (9)
(For convenience we have inserted here a constant $V$ with dimensions of velocity $LT^{-1}$ so that each $z^{i}$ has dimensions of length $L$.)
@@
For $t>t_{1}$ the point moves in a new direction determined by which of the eigenvectors $|\chi_{K}(t_{1})\rangle$ results from application of Born’s Rule (6) to the measurement at $t_{1}$. The possible directions are determined by and associated with the corresponding “velocity vectors” having components $w_{K}^{\,\,\,i}(t_{1})=V\,\beta_{K}^{\,\,\,i}(t_{1})$ with $\beta_{K}^{\,\,\,i}(t_{1})$ as in (3). Note that these directions are orthogonal according to (4). Note also that the probabilities associated with the different directions as given by (4) can be viewed as the moduli squared of generalized direction cosines between the velocity vector immediately before the measurement $v^{i}$ and those possible immediately after the measurement, the $w_{K}^{\,\,\,i}(t_{1})$, since
$|\overline{v^{i}(t_{1})}w_{K}^{\,\,\,i}(t_{1})|^{2}=V^{2}|\overline{\alpha^{i}}\beta_{K}^{\,\,\,i}(t_{1})|^{2}\,.$ (11)
For $t_{0}<t<t_{1}$ the vectors $w_{K}^{\,\,\,i}(t)$ evolve in time as noted above and may be pictured as a group of orthogonal velocity vectors rotating about the state point as it moves along the straight line (10). If the state vector selected after the measurement at $t=t_{1}$ is $|\chi_{{K^{\prime}}}(t_{1})\rangle$ then for $t>t_{1}$ the state point moves with new constant velocity $w_{K^{\prime}}^{\,\,\,i}(t_{1})$, so that
@@
If no observation is made the new state is a mixture of the eigenvectors weighted by the probabilities given by Born’s Rule and the trajectory of the state point belongs to a “fan” of possible trajectories weighted accordingly.
@@
$|\varphi^{\prime}_{i}\rangle={\widehat{U}}|\varphi_{i}\rangle\Rightarrow|\varphi^{\prime}_{i}\rangle=\sum_{j}U_{i}^{\,j}|\varphi_{j}\rangle\,,\qquad\qquad$
@@
and if as before ${\widehat{A}}(t)$, ${\widehat{B}}(t)$dots comprise the set of commuting operators measured at $t=t_{1}$ and the system moves into the state $|\chi\rangle_{K^{\prime}}(t_{1})$ after the measurement, then
$\sum_{j}A_{j}^{\,\,i}(t_{1})\,w_{K^{\prime}}^{\,\,\,j}(t_{1})=a\,w_{K^{\prime}}^{\,\,\,i}(t_{1})\,,$
$\sum_{j}B_{j}^{\,\,i}(t_{1})\,w_{K^{\prime}}^{\,\,\,j}(t_{1})=b\,w_{K^{\prime}}^{\,\,\,i}(t_{1})\,,\dots$ (16)
corresponding to (3) and (8).
Note also that because any state vector $|\psi\rangle$ can be identified with $e^{i\theta}|\psi\rangle$ for every real $\theta$, the space $S$ can be assumed to have the projective property that any two points with coordinates $z^{i}$ and $e^{i\theta}z^{i}$ for all $i=1\,,2\,,\dots$ are to be identified for every real $\theta$.
At this point it is worth emphasizing that the quantum system as described may consist of arbitrarily many interacting particles (or subsystems). Accordingly, the Hilbert space $\cal{H}$ could be the tensor product of many sub-Hilbert spaces.
@@
$|\psi\rangle=\sum_{i\,j\,k}\alpha^{i\,j\,k}|\varphi_{i}\rangle\otimes|\chi_{j}\rangle\otimes|\rho_{k}\rangle$ (17)
in place of (2) and the components of the velocity vector and the coordinates of the state point can accordingly now be labelled $v^{i\,j\,k}$ and $z^{i\,j\,k}(t)$ respectively.
This can be considered a refinement of the description that need not be pursued further for present purposes. What is esential is that the encompassing space $\cal{H}$ has a countable basis.
The extension of the quantum description as described so far can be thought of as an extension of the “matrix mechanics” formulation of quantum dynamics [16] which is as old as quantum mechanics itself and has its origins in the pioneering work of Heisenberg and Born. Observables are represented there by Hermitian matrices, time-dependent in general, and defined as above. All quantum mechanical calculations can now be carried out in terms of the state point and its velocity in $S$. For example the expectation value of an observable $A(t)$ in the state $|\psi\rangle$ between measurements can be expressed as
@@ 3 Hidden variables and quantum measurement
The reader may consider that the behaviour of the state point is analogous to that of a macroscopic object floating down a horizontal stream that forks into two such streams at right angles. The square of the direction cosine between the direction of either fork and that of the original stream may be considered a first estimate of the probability that the object will float down that particular fork. For example if the two forks are at angles of $\pi/6$ and $\pi/3$ with the original stream, with associated direction cosines $\sqrt{3}/2$ and $1/2$ respectively, then the associated probabilities are $3/4$ and $1/4$. The critical difference between this classical behaviour and that following the quantum measurement is that the indeterminacy in the behaviour of the classical object on reaching the fork can in principle be reduced arbitrarily greatly by more refined observation of the system prior to the forks being reached, so that it becomes more certain which fork will be followed. Except in special cases [9] this is not possible in the quantum case.
In short, the classical indeterminacy is arbitrarily reducible in principle whereas in general the quantum indeterminacy is irreducible.
@@ 4 A suggested generalization
It is convenient to address these questions in the context of a natural generalizarion of the dynamical substructure described so far and we now propose that the space underlying quantum dynamics as described above is actually a locally flat subspace of a more general space $S$, a complex Riemannian manifold with associated Hermitian metric tensor
@@
where we have now introduced the summation convention.
In general we can describe the points in $S$ associated with measurements as local singularities or “stagnation points” in $S$ associated with the location in space-time where the measurements take place, being typically the location of measuring devices in a meta-stable state [16] – think of a cloud chamber, a Geiger counter or a photographic plate, for example.
As to the meaning of $S$, we propose that it represents the entropy content (equivalently, the information content) of the physical environment within which the quantum system evolves, including the measuring apparatus. That implies that the structure of $S$ changes in some way when the entropy content of the environment changes. For example when a photographic plate is exposed during a quantum measurement it is clear that the entropy of the neighbouring environment increases abruptly and the associated singularity in $S$ disappears. More generally “measurements” may simply refer to interactions between the system and its environment at singular points in $S$. In the absence of an observer the system moves after each such interaction into a more and more complicated mixed state. Whether such interactions provide the only source for changes in the structure of $S$ and its curvature remains an open question. Perhaps further analysis will lead to the formulation of an equation similar to Einstein’s equation [23], relating change in the curvature of $S$ to an analogue of the energy-momentun tensor.
Turning to the behaviour of the state point in $S$ between measurements, straight line motion as in (10) is naturally generalized to motion along a geodesic between the locations at $P$ and $Q$ say, of quantum measurements, so minimizing the distance travelled. This may be regarded as a generalization to $S$ of the principle of least action that leads to geodesic motion of a mass point (a “test particle”) in space-time [23] and leads here to a variational condition in the familiar form [15]
[…]
@@
$g_{i\,j}\,\frac{dp^{j}}{du}-\frac{\partial g_{k\,l}}{\partial z^{i}}\,\bar{p}^{k}p^{l}+\frac{\partial g_{j\,i}}{\partial z^{k}}\bar{p}^{j}p^{k}+\frac{\partial g_{i\,j}}{\partial z^{k}}p^{j}p^{k}+\quad{\rm{c.\,c.}}=0\,.$ (22)
@@
$\frac{\delta p^{m}}{\delta u}=g^{m\,i}\left(g_{i\,j}\,\frac{dp^{j}}{du}-\frac{\partial g_{k\,l}}{\partial z^{i}}\,\bar{p}^{k}p^{l}+\frac{\partial g_{j\,i}}{\partial z^{k}}\bar{p}^{j}p^{k}+\frac{\partial g_{i\,j}}{\partial z^{k}}p^{j}p^{k}\right)$
$=\frac{dp^{m}}{du}-g^{m\,i}\left(\frac{\partial g_{k\,l}}{\partial z^{i}}\,\bar{p}^{k}p^{l}-\frac{\partial g_{j\,i}}{\partial z^{k}}\bar{p}^{j}p^{k}-\frac{\partial g_{i\,j}}{\partial z^{k}}p^{j}p^{k}\right)$ (24)
@@
The vector (with components) $v^{i}$ corresponding to the the state vector $|\psi\rangle$ is now to be considered as parallel transported along the geodesic traced by $z^{i}(t)$, where $u$ is now replaced by elapsed time $t$ along that geodesic and $p^{i}(u)$ ($=dz^{i}/du$) is replaced by $v^{i}(t)$ ($=dz^{i}/dt$). The state vector is no longer a constant vector in general between measurements and $v^{i}$ has vanishing absolute derivative along the geodesic as defined from (24) by
$\frac{\delta v^{m}}{\delta t}=\frac{dv^{m}}{dt}-g^{m\,i}\,\left(\frac{\partial g_{k\,l}}{\partial z^{i}}\,\bar{v}^{k}v^{l}-\frac{\partial g_{j\,i}}{\partial z^{k}}\bar{v}^{j}v^{k}-\frac{\partial g_{i\,j}}{\partial z^{k}}v^{j}v^{k}\right)=0\,.$ (25)
The Hilbert space $\cal{H}$ is now taken to be the tangent space to the geodesic followed by $z^{i}(t)$, obtained by integrating (24). Vectors in $\cal{H}$, including the quantum state vector $|\psi\rangle$, are parallel transported along the geodesic, preserving lengths and orthogonality relations.
It is natural to assume further that the Hamiltonian operator ${\widehat{H}}$ has vanishing absolute derivative along the geodesic, obtained by regarding its matrix representation $H_{j}^{\,\,i}$ as a mixed tensor and generalizing (25) accordingly [15], while the governing differential equation for other operators representing time-dependent observables includes an extra term generalizing (1).
@@
The differences between the simple substructure described in the preceding sections and the generalized substructure of the quantum dynamics speculated upon here may have implications for the outcome of the measurement at $t=t_{1}$ and for the quantum dynamics more generally, not least because the state vector is no longer constant between measurements in this generalized Heisenberg picure. More analysis is necessary to determine the nature of these implications.

### v5

@@ Abstract
The description of a closed quantum system is extended with the identification of an underlying substructure enabling an expanded formulation of dynamics in the Heisenberg picture. Between measurements a “state point” moves in an underlying multi-dimensional complex projective space with constant velocity determined by the quantum state vector. Following a measurement the point changes direction and moves with new constant velocity along one of several possible new orthogonal paths with probabilities determined by the Born Interpretation of the state vector. From this previously hidden substructure a new picture of quantum dynamics and quantum measurements emerges without violating existing no-go theorems regarding hidden variables. A natural generalisation to a Riemannian substructure is proposed, determined by the entropy of the background environment. This leads to a suggested interaction between the substructure of quantum dynamics and the background gravitational field.
email: ajb@maths.uq.edu.au
Keywords: quantum dynamics; geometric substructure; quantum measurements; quantum mechanics and gravitation.
ORCID: 0000-0001-5181-4483
@@ 1 Introduction
[…]
@@
Here $^{\widehat{H}}$ is the Hamiltonian operator. (Only closed systems for which $^{\widehat{H}}$ is time-independent are considered at this stage.)
But in another and very different type of dynamical process that is assumed to reflect the interaction of the system with the measuring apparatus and the observer, a maximal set of commuting observables is chosen for measurement and the state vector is subsequently observed to move into a common eigenvector of that set with a probability determined by Born’s Rule [9, 14, 15]. (Complications that arise with observables having partially continuous spectra are set aside here.)
Suppose that at time $t_{0}$ a measurement has been made of such a set of commuting observables and $|\psi\rangle$ has been observed subsequently in one of their common eigenvectors. The state vector remains constant for $t>t_{0}$ until another measurement takes place. We may expand it during such a time interval in terms of some chosen reference basis, a complete set of orthonormal vectors $|\varphi_{i}\rangle$, $i=1\,,2\,\dots\,$, possibly infinite in number, with constant complex expansion coefficients $\alpha^{i}$. Then
@@
[…]
@@
(In the special case when $^{\widehat{H}}$ is one of the chosen set then all the observables in the set are constants of the motion and their common eigenvectors $|\chi_{K}\rangle$ and all their components $\beta_{K}^{\,\,\,i}$ are constant.)
@@
${\widehat{B}}(t_{1})\,|\psi\rangle={\widehat{B}}(t_{1})|\chi_{K^{\prime}}(t_{1})\rangle=b|\chi_{K^{\prime}}(t_{1})\rangle=b\,|\psi\rangle\,$$\quad\dots\$(8)
@@ 2 Identifying a substructure
The reader may observe that the dynamical process described above is strongly reminiscent of the behaviour of a free particle travelling with constant velocity between impulsive forces applied at $t_{0}$ and $t_{1}$. This suggests the association of the state vector $|\psi\rangle$ with the velocity vector of a point moving in a hitherto unidentified underlying space. Accordingly the $V\,\alpha^{i}$ are now identified with the components $v^{i}$ of the constant velocity vector of a “state point” moving in an underlying space $S$ say, with complex coordinates $z^{i}$ so that
$v^{i}=\frac{dz^{i}}{dt}\,,\qquad i=1\,,2\,$$\dots$ (9)
(For convenience a constant $V$ with dimensions of velocity $LT^{-1}$ has been inserted here so that each $z^{i}$ has dimensions of length $L$.) Note from (2) that
$\sum_{i}\overline{v^{i}}v^{i}=V^{2}\,.$ (10)
@@
For $t>t_{1}$ the point moves in a new direction determined by which of the eigenvectors $|\chi_{K}(t_{1})\rangle$ results from application of Born’s Rule (6) to the measurement at $t_{1}$. The possible directions are determined by and associated with the corresponding “velocity vectors” having components $w_{K}^{\,\,\,i}(t_{1})=V\,\beta_{K}^{\,\,\,i}(t_{1})$ with $\beta_{K}^{\,\,\,i}(t_{1})$ as in (3). Note that these directions are orthogonal according to (4) and that each velocity vector is normalized as in (10). Note also that the probabilities associated with the different directions as given by (6) can be viewed as the moduli squared of generalized direction cosines between the velocity vector immediately before the measurement $v^{i}$ and those possible immediately after the measurement, the $w_{K}^{\,\,\,i}(t_{1})$, since
$\sum_{i}|\overline{v^{i}(t_{1})}w_{K}^{\,\,\,i}(t_{1})|^{2}=V^{2}\sum_{i}|\overline{\alpha^{i}}\beta_{K}^{\,\,\,i}(t_{1})|^{2}=V^{2}P_{K}\,.$ (12)
For $t_{0}<t<t_{1}$ the vectors $w_{K}^{\,\,\,i}(t)$ evolve in time as noted above and may be pictured as a set of orthogonal velocity vectors rotating unitarily about the state point as it moves along the straight line (11). If the state vector observed following the measurement at $t=t_{1}$ is $|\chi_{{K^{\prime}}}(t_{1})\rangle$ then for $t>t_{1}$ the state point moves with new constant velocity $w_{K^{\prime}}^{\,\,\,i}(t_{1})$, so that
@@
For $t<t_{1}$ it is indeterminate which of the evolving velocity vectors $w_{K}^{\,\,\,i}(t)$ will result from the measurement at $t_{1}$. If no observation is made the new state for $t>t_{1}$ is a mixture of the eigenvectors weighted by the probabilities given by Born’s Rule and the trajectory of the state point belongs to a “fan” of possible trajectories weighted accordingly.
@@
$|\varphi^{\prime}_{i}\rangle={\widehat{U}}|\varphi_{i}\rangle\Rightarrow|\varphi^{\prime}_{i}\rangle=\sum_{j}U_{i}^{\,j}|\varphi_{j}\rangle\,$$\qquad\qquad$
@@
In addition to $^{\widehat{U}}$, the action of all linear operators on the Hilbert space of state vectors can be extended to action as matrices on the state point and its velocity in $S$, for example
and if as before ${\widehat{A}}(t)$, ${\widehat{B}}(t)$ $\dots$ comprise the set of commuting operators measured at $t=t_{1}$ and the system moves into the state $|\chi\rangle_{K^{\prime}}(t_{1})$ after the measurement, then
$\sum_{j}A_{j}^{\,\,i}(t_{1})\,w_{K^{\prime}}^{\,\,\,j}(t_{1})=a\,w_{K^{\prime}}^{\,\,\,i}(t_{1})\,$,
$\sum_{j}B_{j}^{\,\,i}(t_{1})\,w_{K^{\prime}}^{\,\,\,j}(t_{1})=b\,w_{K^{\prime}}^{\,\,\,i}(t_{1})\,$$\dots$ (17)
corresponding to (3) and (8).
@@
[…]
Unlike the closed system on which attention has been focussed so far some of these extra systems will be open and interacting continuously or intermittently, unitarily or non-unitarily with their environment rather than at isolated points associated with measurements. They enter increasingly complicated mixed states as time passes [16] - [18].
(These extensions of the description to include multiple quantum systems, whether interacting or not and whether open or not, are not pursued here. Furthermore the complications arising from the introduction of quantum fields are not considered.)
The generalised description of one closed quantum system as described above can be thought of as an extension of the “matrix mechanics” formulation of quantum dynamics [19] which is as old as quantum mechanics itself and has its origins in the pioneering work of Heisenberg and Born. Observables are represented there by Hermitian matrices, time-dependent in general, and defined as above. All quantum mechanical calculations can now be carried out in terms of the state point and its velocity. For example the expectation value of an observable $A(t)$ in the state $|\psi\rangle$ between measurements can be expressed as
@@ 3 Hidden variables and quantum measurement
Before the measurement the Shannon - von Neumann entropy of the system has the value
$-\rho\log(\rho)=0$ (19)
where $\rho$ denotes the density matrix [9], which has just the one eigenvalue $1$ when the system is in a pure state $|\psi\rangle$ and
$\rho=|\psi\rangle\langle\psi|\,.$ (20)
Then the state point has a definite trajectory as in (11). After the measurement at $t=t_{1}$
$\rho=\sum_{K}P_{K}|\chi_{K}(t_{1})\rangle\langle\chi_{K}(t_{1})|$ (21)
and the entropy takes the larger value
$\sum_{K}P_{K}\log(1/P_{K})\,.$ (22)
Here the $P_{K}$ are the Born probabilities of the trajectories in the fan as in (6). The entropy decreases to 0 again after the observation of which of those trajectories the state point now follows, with the system again in a pure state.
Only simple measurements are considered here as described in [9, 14, 15] and by many others since. More sophisticated descriptions of the measurement process and more generally of a quantum system’s possible interactions with its environment have been developed over the years – see in particular [16, 17, 18, 20] – together with more sophisticated measures of the system’s entropy [27].
The reader may consider that the behaviour of the state point before and after a measurement is analogous to that of a macroscopic object floating down a horizontal stream that forks into two such streams at right angles. The square of the direction cosine between the direction of either fork and that of the original stream may be considered a first estimate of the probability that the object will float down that particular fork. For example if the two forks are at angles of $\pi/6$ and $\pi/3$ with the original stream, with associated direction cosines $\sqrt{3}/2$ and $1/2$ respectively, then the associated probabilities are $3/4$ and $1/4$.
The critical difference between this classical behaviour and that associated with the quantum measurement is that the indeterminacy in the behaviour of the classical object prior to reaching the fork can in principle be reduced arbitrarily greatly by more refined observation of the system so that it becomes more certain which fork will be followed. Except in special cases [10] this is not possible in the quantum case. In short, the classical indeterminacy prior to the fork is arbitrarily reducible in principle whereas in general the quantum indeterminacy prior to the measurement – associated with the set of velocity vectors rotating unitarily about the trajectory of the state point – is irreducible.
@@ 4 A suggested generalization
It is convenient to address these questions in the context of a natural generalizarion of the dynamical substructure described so far and it is now proposed that the space underlying quantum dynamics as described above is actually a locally flat subspace of a more general space $S$, a complex Riemannian manifold with associated Hermitian metric tensor
@@
where the summation convention has now been introduced.
In general the points in $S$ associated with measurements of a closed quantum system can be described as local singularities or “stagnation points” in $S$ associated with the location in space-time where the measurements take place, being typically the location of measuring devices in a meta-stable state [19] – think of a cloud chamber, a Geiger counter or a photographic plate, for example.
As to the meaning of $S$, it is proposed that it represents the entropy content (equivalently, the information content) of the physical environment within which quantum systems evolve, including any measuring devices. This implies that the structure of $S$ changes when the entropy content of the environment changes. For example when a photographic plate is exposed during a quantum measurement it is clear that the entropy of the neighbouring environment increases abruptly as the associated singularity in $S$ disappears. More generally “measurements” may simply refer to interactions between quantum systems and their environment whether continuous or stochastic, at singular points in $S$ or non-locally. In the absence of observations quantum systems move after such interactions into more and more complicated mixed states with higher entropy. It is suggested that while such interactions increase the entropy of the quantum systems involved and their environment, they also provide the (only) source for changes in the structure of $S$.
One possible description of that structure would be provided by supposing that it is a Kähler manifold [7, 28] with real potential $B(\bar{z},z)$ such that
$g_{i\,j}(\bar{z}\,,z)=\frac{\partial^{2}B(\bar{z},z)}{\partial{\bar{z}}^{i}\partial{z}^{j}}\,$, (25)
and by supposing further that $B$ is a measure of the entropy/information content of the environment in which quantum systems evolve. It would remain to determine a mathematical description of how changes in $B$ arise from interactions of those systems with their environment. Note that (24) and (25) would imply
$ds^{2}=\frac{\partial^{2}B(\bar{z},z)}{\partial{\bar{z}}^{i}\partial{z}^{j}}\,d{\bar{z}}^{i}\,dz^{j}\qquad$
$\rm{and}\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad$
$\frac{\partial g_{i\,j}}{\partial z^{k}}=\frac{\partial g_{i\,k}}{\partial z^{j}}\,,\quad\frac{\partial g_{i\,j}}{\partial\bar{z}^{k}}=\frac{\partial g_{i\,k}}{\partial\bar{z}^{j}}\,.$ (26)
Bearing (25) in mind as a possibility, while returning to the general discussion of the behaviour of the state point of a closed system between measurements, it is clear that straight line motion as in (11) is naturally generalized to motion along a geodesic between the locations at $P$ and $Q$ say, of quantum measurements, so minimizing the distance travelled. This may be regarded as an analogue of the principle of least action that leads to geodesic motion of a mass point (a “test particle”) in space-time [30] and leads here to a variational condition in the familiar form [29]
$0=\delta\,\int_{P}^{Q}\,ds=\int_{u_{P}}^{u_{Q}}\left(g_{i\,j}\bar{p}^{i}p^{j}\right)^{1/2}\,du\,,\quad\bar{p}^{i}=\frac{d{\bar{z}}^{i}}{du}\,,\quad p^{j}=\frac{dz^{j}}{du}\,$, (27)
@@
$g_{i\,j}\,\frac{dp^{j}}{du}-\frac{\partial g_{l\,k}}{\partial{\bar{z}}^{i}}\,\bar{p}^{l}p^{k}+\frac{\partial g_{i\,j}}{\partial{\bar{z}}^{k}}\bar{p}^{k}p^{j}+\frac{\partial g_{i\,j}}{\partial z^{k}}p^{j}p^{k}+\quad{\rm{c.\,c.}}=0\,.$ (28)
@@
$\frac{\delta p^{m}}{\delta u}=g^{m\,i}\left(g_{i\,j}\,\frac{dp^{j}}{du}-\frac{\partial g_{l\,k}}{\partial{\bar{z}}^{i}}\,\bar{p}^{l}p^{k}+\frac{\partial g_{i\,j}}{\partial{\bar{z}}^{k}}\bar{p}^{k}p^{j}+\frac{\partial g_{i\,j}}{\partial z^{k}}p^{j}p^{k}\right)$
$=\frac{dp^{m}}{du}-g^{m\,i}\left(\frac{\partial g_{l\,k}}{\partial{\bar{z}}^{i}}\,\bar{p}^{l}p^{k}-\frac{\partial g_{i\,j}}{\partial{\bar{z}}^{k}}\bar{p}^{k}p^{j}-\frac{\partial g_{i\,j}}{\partial z^{k}}p^{j}p^{k}\right)$ (30)
@@
The vector (with components) $v^{i}$ corresponding to the the state vector $|\psi\rangle$ is now to be considered as parallel transported along the geodesic traced by $z^{i}(t)$, where $u$ is now replaced by elapsed time $t$ along that geodesic and $p^{i}(u)$ ($dz^{i}/du$) is replaced by $v^{i}(t)$ ($dz^{i}/dt$). The state vector and corresponding velocity vector $v^{i}$ are no longer constant between measurements. Instead $v^{i}$ has vanishing absolute derivative along the geodesic as defined from (30), so that
$0=\frac{\delta v^{m}}{\delta t}=\frac{dv^{m}}{dt}-g^{m\,i}\,\left(\frac{\partial g_{l\,k}}{\partial{\bar{z}}^{i}}\,\bar{v}^{l}v^{k}-\frac{\partial g_{i\,j}}{\partial{\bar{z}}^{k}}\bar{v}^{k}v^{j}-\frac{\partial g_{i\,j}}{\partial z^{k}}v^{j}v^{k}\right)\,.$ (31)
Note that all of (23)-(27) and (29)-(31) are consistent with the projective condition introduced in Sec. 2 which identifies $z^{i}$ with $e^{i\theta}z^{i}$ for all values of $i$ and every real $\theta$ and which now carries over to the generalisation of $S$ to a Riemannian manifold, whether Kähler or not.
[…]
It is natural to assume further that the Hamiltonian operator $^{\widehat{H}}$ has vanishing absolute derivative along the geodesic, obtained by regarding its matrix representation $H_{j}^{\,\,i}$ as a mixed tensor and generalizing (31) accordingly [29], while the governing differential equation for other operators representing time-dependent observables includes an extra term generalizing (1).
@@
The differences between the simple substructure described in the preceding sections and the generalized substructure of the quantum dynamics proposed here can be expected to have implications for the outcome of quantum measurements and for quantum dynamics more generally, not least because of possible effects of the curvature of $S$, in particular on the evolution of the state vector in accordance with (31). More analysis is necessary to determine the nature of such implications.